Which of the following algebraic expressions are rational: a.) x^2-3x+4x=8/4, b.) 1/(3x) + 4x/6 =1, c.) 5/x +x/9=2/4, d.) x^2 -3x+4x=0, e.) 3+ (4x/6)=1
1
Expert's answer
2020-10-20T18:53:18-0400

 A rational expression is a fraction in which the denominator and/or the numerator are polynomials.


Let us analyse each expressions with respect to the definition given above.


a). x²-3x+4x =8/4

This can be written as "\frac{x²-3x+4x}{1}" = "\frac{8}{4}"

Since the equation can be written in this form, that implies, it is rational.


b). 1/(3x) + 4x/6 =1

Let us make the L.H.S. a single fraction

"\frac{1}{3x} + \frac{4x}{6} = 1"

Find the L.C.M.

"\frac{2+4x²}{6x} = 1"

Since the equation 1/(3x) + 4x/6 =1 can be written as "\frac{2+4x²}{6x} = \frac{1}{1}"

That implies it is rational.


c).5/x +x/9=2/4

Let us make the L.H.S. a single fraction.

"\frac{5}{x} + \frac{x}{9} = \frac{2}{4}"

Find the L.C.M.

"\frac{45+x²}{9x} = \frac{2}{4}"


Since the equation 5/x +x/9=2/4 can be written as "\frac{45+x²}{9x} = \frac{2}{4}"

That implies it is rational


d).x² -3x+4x=0

This can be written as "\frac{x²-3x+4x}{1}=\frac{0}{1}"

Since the equation can be written in this form, that implies it is rational.


e).3+ (4x/6)=1

Let us make the L.H.S. a single fraction

"\frac{3}{1} + \frac{4x}{6} = \frac{1}{1}"

Find the L.C.M.

"\frac{18+4x}{6} = \frac{1}{1}"

Since the equation 3+ (4x/6)=1 can be written as "\frac{18+4x}{6} = \frac{1}{1}"

That implies it is rational.